OpenStudy (mitchi35):

I realize that this is probably really simple. But I just don't understand how to do this. Help please? :) 1. Using your calculator, complete the table below for the function y = cos x. Find the corresponding y-values rounded to the nearest thousandth. I just need help with the fist one.

OpenStudy (mitchi35):

On the table these are the x values(they are all degrees)and I need to figure out the y value: x:0 30 60 90 120 150 180 210 240 270 300 330 360 y: Next table has the same ones but they are all negatives(still degrees) and I need the y value x: 0 -30 -60 -120 -150 -180 -210 -270 -300 -330 -360 y:

OpenStudy (shikamaru11):

well what kinda relation is given there is it mentioned???? in the table.....

OpenStudy (mitchi35):

They both have the same numbers but one is all positive, and the other are negatives?

OpenStudy (shikamaru11):

so the x and should be inverse of each other ???/ what's your mind says...

OpenStudy (shikamaru11):

means to say X = 1/Y or Y=1/X ??? is that

OpenStudy (mitchi35):

Oh. Then no I guess.

OpenStudy (shikamaru11):

so do you know the answer??? :)

OpenStudy (mitchi35):

No I don't.... :(

OpenStudy (shikamaru11):

ok wait a minute

OpenStudy (shikamaru11):

X = -Y for the 1st and -X = +Y :) did you get it now

OpenStudy (mitchi35):

Oh! okay. But how do I figure out the rest of the points? And sorry, I had to run somewhere real quick. :)

OpenStudy (shikamaru11):

oh It's ok ma'm well guess it easily simple have you seen examples related to it??

OpenStudy (mitchi35):

I have seen some of them I just did't understand them that well by myself. They have some of them in the lesson plan.

OpenStudy (shikamaru11):

@Mitchi35 well sorry for being late so now btw what's your grade ma'm??

OpenStudy (mitchi35):

I have a B-. So if I don't get this, this can bring my grade down to a C. It's worth 50 points. :(

OpenStudy (mitchi35):

Why do you ask?

OpenStudy (shikamaru11):

oh not this means your class @Mitchi35 :) well don't worry if you're taking pain then you'll gain too

OpenStudy (mitchi35):

Okay thanks. But lol can you help me with this orrrrrrr?...

Directrix (directrix):

Use the calculator here. http://web2.0calc.com/ @Mitchi35